# BlockFUL: Enabling Unlearning in Blockchained Federated Learning

### v2

@@ III-E Design Goal
IV Proposed Dual-Chain Unlearning Architecture
In this section, we design a new dual-chain structure in the BlockFUL framework, where an archive chain is utilized to record historical models and a live chain is employed to share the up-to-date FL models. We design the block structure in the live chain using primitives, such as the CH algorithm, to provide redactable operations on the data stored on the live chain to enable the process of the blockchain unlearning stage.
IV-A Transactional Models in the Dual-Chain Structure
This process records models in the blockchain. In our dual-chain structure, each transaction represents a model, and each block contains multiple transactions. These transactions are uploaded to the archive and live chains separately, through unified consensus, thereby reducing the frequency of the consensus process in the entire blockchain network. Models are stored in the Inter Planetary File System (IPFS) after undergoing consensus and validation. Note that the models of the archive chain and the live chain are stored in different IPFS instances. Fig. 3 shows the proposed dual-chain structure.
Archive Chain. The block data records generated by local participation in FL are uploaded to the archive chain. The archive chain does not grant access permission to FL users; only relevant auditing parties can view model records in the archive chain when necessary. This approach provides traceability for all model uploading activities and ensures the integrity of the models. The archive chain needs to record all complete historical models before and after unlearning. These historical models do not need updating. In this case, a smart contract manages historical models in a list format.
Live Chain. The transaction records generated by local training models are uploaded to the live chain. The live chain provides model sharing and includes the latest model versions. Model updates are executed by the committee. The models are only updated on the live chain when all committee members reach a unanimous consensus. The structure of the live chain differs from that of the archive chain. Since the live chain only records the latest state of the models, smart contracts need to operate with models as transaction entries.
@@ IV-A Transactional Models in the Dual-Chain Structure
  • $CH_{curr}\left(h\right)$: Considering a model update task, it functions as the hash value for the current block. This role ensures that any changes to the model version resulting from updates, do not affect the validity of the current block.
@@
  • $version$: This field is updated with the corresponding version number for each update task, maintaining consistent versions in the global state.
There are also the common block body’s Merkle root $MR(h)$ and the current block generation time $timestamp$ in the block header.
Block body. A block body contains multiple redactable transactions uniformly composed in a Merkle tree structure and stores model-related values. A redactable transaction contains the following fields, where, we take transaction $Tx_{i,j}$ as an example, i.e., the $j$-th transaction in the body of block $i$.
@@
The design of this block contributes to efficient and secure redactability in the blockchain. Due to the use of CH, transaction updates do not alter the block body’s Merkle root. However, the version field in the block header may cause collaborative updating issues with the block header. To ensure the consistency of transaction versions, we introduce the $version$ field in the block body. As a result, the validity of the link between blocks is maintained, and computational overhead is minimized.
Moreover, $CH\left(i,j\right)$ is a field in the block that needs verification for the block body. In this way, even if $URI_{w}\left(i,j\right)$ and $v(i,j)$ are updated, $CH(i,j)$ in the block body remains unchanged. Thus, $MR(h)$ in the block header also remains unchanged. These conditions satisfy the immutability of $CH_{curr}\left(h\right)$ and $CH(i,j)$. A valid block with an immutable block header can still be verified.
On the other hand, this redactability does not compromise the tamper-resistance of the blockchain. The redactability of blocks is guaranteed by $CH_{sk}$ (trapdoor), an indispensable input for CH updates [27]. Meanwhile, to prevent malicious committee members from deliberately exposing outdated transaction information, once a transaction becomes outdated, it is not stored, but directly discarded instead.
IV-B Model Unlearning in the Live Chain
@@
The parameter update equation related to $CH$, as specified in (5), ensures that the original $CH$ value remains valid after information updates. The committee inputs the original payload $\mathcal{T}(i,j)$, which includes the model $URI_{w}(i,j)$ before the update, along with the new payload $\mathcal{T}^{\prime}(i,j)$ containing the updated model $URI^{\prime}_{w}(i,j)$ and the unchanged reference list $R_{Tx}(i,j)$, the random number $v(i,j)$ before the update, and the corresponding private key $CH{sk}$. It outputs the updated random number $v^{\prime}(i,j)$, as follows:
$\begin{split}CH\ update\left(\mathcal{T}(i,j),\mathcal{T}^{\prime}(i,j),v(i,j),CH_{sk}\right)\rightarrow\\ v^{\prime}(i,j).\end{split}$ (5)
@@
After the transaction update is completed, the block header also needs to be updated. This can be achieved by updating the version field, following a similar process to the transaction update. Once the committee’s consensus nodes complete the transaction and block updates, the committee sends the updates to the other consensus nodes to complete the entire consensus process.
V Blockchained Federated Unlearning Paradigms
@@
V-A Paradigm 1 - Parallel Unlearning
In this paper, “parallel unlearning” refers to the committee members collectively reaching a single consensus for updating all to-be-unlearned models within an unlearning task. After updating all models and reaching a unified consensus, the results are recorded on the chain simultaneously. To illustrate the parallel unlearning paradigm, we employ the gradient ascent to demonstrate this process. We introduce a new gradient passing function that manipulates the ascending gradients for inherited models within the DAG using the ascending gradients from the starting models. Then, the inherited models can be updated in parallel using the generated ascending gradients. In the function design, we prioritize the unlearning results across all inherited models by highlighting the inheritance relationships from the starting models to the inherited models.
@@ V-A Paradigm 1 - Parallel Unlearning
$w^{\prime}_{y}=\sum_{p_{i}\in P}\frac{1}{\prod_{w_{j}\text{ on }p_{i}}N^{R_{j}}}\times\alpha\nabla\theta_{s}+w_{y},$ (7)
where $P$ denotes the set of paths from $w_{s}$ to $w_{y}$, $P=\left\{p_{1},\ldots,p_{l}\right\}$, “$w_{j}\text{ on }p_{i}$” denotes traversing each model node on path $p_{i}$ except $w_{s}$, $N^{R_{j}}$ denotes the number of models inherited by model node $w_{j}$ passing through path $p_{i}$, and $\alpha$ denotes discount factor. Equation (7) handles model inheritance in FL, enabling the update of all relevant inherited models.
Update with multiple starting models. In this case, a user participating in a BlockFUL task contributes multiple models, and these models may have different gradients. During updates, these models serve as the starting models and have varying degrees of influence on the inherited models (i.e., the update of the inherited models may involve the computation of one or more starting model gradients). In other words, the inherited models may be affected by scenarios, for example, the same gradient but different paths, or different gradients and different paths.
@@
$w^{\prime}_{y}=\!\sum_{\begin{gathered}\nabla\theta^{k}_{s}\!\in\!\nabla\theta\\ P^{\nabla\theta^{k}_{s}}\!\in\!\mathbb{P},\end{gathered}}\left(\sum_{p^{\nabla\theta^{k}_{s}}_{i}\in P^{\nabla\theta^{k}_{s}}}\frac{1}{\prod_{w_{j}\text{ on}p_{i}}N^{R_{j}}}\!\times\!\alpha\nabla\theta^{k}_{s}\right)\!+w_{y},$ (8)
where $\nabla\theta$ denotes the set of multiple model gradients for a user, and $\nabla\theta=\left\{\nabla\theta^{1}_{s},\nabla\theta^{2}_{s},\ldots,\nabla\theta^{h}_{s}\right\}$; $p^{\nabla\theta^{k}_{s}}_{i}$ denotes the $i$-th path with gradient $\nabla\theta^{k}_{s}$; and $P^{\nabla\theta^{k}_{s}}$ denotes the set of paths with gradient $\nabla\theta^{k}_{s}$. Additionally, $\mathbb{P}$ is the set of paths with all gradients, i.e., =$\mathbb{P}=\left\{P^{\nabla\theta^{1}_{s}},\ldots,P^{\nabla\theta^{h}_{s}}\right\}$.
@@
Next, we prove the depths required from the starting model to its inherited model to cease updating are bounded. By denoting $\nabla\theta$ as $\alpha\nabla\theta_{s}$, and considering $D=\{d_{1},d_{2},\ldots,d_{l}\}$ with $d_{i}$ being the depth of path $p_{i}$, we have the following theorem.
@@
First, from (7), it readily follows that
@@
$\begin{split}&\sum_{p_{i}\in P}\frac{|\nabla\theta|}{\prod_{w_{j}\text{ on }p_{i}}N^{R_{j}}}\leqslant\sum_{d_{i}\in D}\frac{|\nabla\theta|}{2^{d_{i}}}\leqslant\varepsilon\\ &\Longrightarrow\sum_{d_{i}\in D}\frac{1}{2^{d_{i}}}\leqslant\frac{\varepsilon}{|\nabla\theta|}.\end{split}$ (15)
If $\frac{1}{\prod_{w_{j}\text{ on }p_{i}}N^{R_{j}}}=\frac{1}{2^{d_{i}}}$ at this point, the value of $d_{i}$ can be obtained as
@@
According to (8), assuming model $w_{y}$ is influenced by multiple distinct gradient model updates, we can represent $B^{k}_{s}$ the gradient ascent required for $w_{y}$ under the influence of gradient $\nabla\theta^{k}_{s}$. Here, $B^{k}_{s}=\sum_{p^{\nabla\theta^{k}_{s}}_{i}\in P^{\nabla\theta^{k}_{s}}}\frac{1}{\prod\nolimits{{w_{j}\text{ on }p_{i}}}N^{R_{j}}}\times\alpha\nabla\theta^{k}_{s}$. Since multiple starting models are updated simultaneously, the gradient ascent process accumulates individual gradient ascent values. In other words, $B=B^{1}_{s}+B^{2}_{s}+\cdots+B^{h}_{s}$. According to Theorem 1, the inherited models influenced by the starting models eventually stop updating when a certain depth is reached. Therefore, as $\prod\nolimits_{w_{j}\text{ on }p_{i}}N^{R_{j}}$ increases, $B^{k}_{s}\leqslant\varepsilon$, further resulting in $B\leqslant\varepsilon$.
Parallel unlearning in BlockFUL. In Algorithm 1, the users broadcast the updated model message to the network. Upon receiving the message, the committee processes the starting updated model set $W$ and its inherited model set $N(W)$. According to (8), the committee performs gradient ascent updates on the inherited models. Similarly, these updates need to be executed on the blockchain. Since the reference list $R_{Tx}$ in each transaction provides reference relations for all transactions in the BlockFUL task. With the assistance of $CH_{sk}$, each committee member can overwrite all transactions and versions to be updated in the blockchain. In this parallel unlearning, we only need to perform one consensus operation to update FL and the blockchain. Based on all these previous updates, a unified hash value can be computed. This hash value is used to execute the consensus process, which helps prevent malicious committee members from making false updates.
@@ 6       Send model_update message($w^{\prime}_{i}$);
9             $w^{\prime}_{j}$leftarrow Committee.Update($w_{j}$, $\nabla\theta$, $\alpha$);
@@ 12            $\mathcal{T}^{\prime}(i,j)$, $v_{i,j},CH_{sk}$);
13             $R^{\prime}_{curr}$leftarrow Committee.Overwrite($version$,
14            $new$ $version$, $R_{curr},CH_{sk}$);
15
@@ 14            $new$ $version$, $R_{curr},CH_{sk}$);
19
20while Consensus do
21       Committee.Verify$(Hash(W^{\prime},N(W^{\prime})),v^{\prime}_{i,j},R^{\prime}_{curr})$;
22       if $Hash(W^{\prime},N(W^{\prime}),v^{\prime}_{i,j},R^{\prime}_{curr})$ $is$ $invalid$ then
@@ 24             Update the corresponding parameters in IPFS;
Algorithm 1 Parallel unlearning (gradient ascent)
V-B Paradigm 2 - Sequential Unlearning
In this paper, “sequential unlearning” refers to a process where committee members execute a consensus process for each individual model update, i.e., a single DAG node, within an unlearning task. After updating one model and reaching a consensus, the results are recorded on the chain before moving on to the next model update. In the sequential unlearning paradigm, we employ the re-training method to demonstrate this process. Here, the same user model settings are used as in gradient ascent. The re-training starts from the updated model $w^{\prime}_{s}$ and goes through all paths to the model $w_{y}$.
V-B1 Re-aggregation
@@ V-B2 Re-training
Sequential unlearning in BlockFUL. The steps for model updating using sequential unlearning, i.e., the re-training method, are shown in Algorithm 2. In this process, the client first checks whether it is the starting model. If it is not the starting model, the client performs the operations of re-aggregation and re-training of its own model locally. Then, the client broadcasts the model update message to the network. The committee utilizes $CH_{sk}$ in the blockchain to perform the same updating operations as described in Algorithm 1. Subsequently, consensus verification is conducted. Upon reaching a consensus, the committee updates the blocks and parameters accordingly in the IPFS. Finally, the committee sends model update messages to the next-hop clients for updating the one-hop inherited models $N^{1}(W)$, two-hop inherited models $N^{2}(W)$, and so on and so forth, where all these inherited models of different hops belong to $N(W)$.
@@
2Define All functions inherited from Algo. 1;

### v3

@@ IV-E Design Goal
V Dual-Chain Unlearning Architecture
In this section, we design a new dual-chain structure in the BlockFUL framework, where an archive chain is utilized to record historical models and a live chain shares the up-to-date FL models. This offers unlearning capabilities while ensuring traceability and integrity. We design the block structure in the live chain using primitives, e.g., the CH algorithm, to provide redactable operations on the data stored on the live chain to enable the process of the blockchain unlearning stage.
@@ V Dual-Chain Unlearning Architecture
V-A Transactional Models in the Dual-Chain Structure
This process records models in the blockchain using a dual-chain structure; see Fig. 3. Each transaction represents a model, with multiple transactions per block. Transactions are uploaded to archive and live chains via unified consensus, minimizing network consensus frequency. After validation, models are stored in decentralized storage, e.g., Inter Planetary File System (IPFS), with separate instances for each chain.
Archive Chain. The block data records generated by local participants in FL are uploaded to the archive chain. The archive chain does not grant access permission to FL users; only relevant auditing parties can view model records in the archive chain. This approach provides traceability for all model uploading activities and ensures the integrity of the models. The archive chain needs to record all historical models before and after unlearning. These historical models do not need updating, and are managed by a smart contract in a list format.
Live Chain. The transaction records generated by local training models are uploaded to the live chain. The live chain provides model sharing and includes the latest model versions. Model updates are executed by the committee. The models are only updated on the live chain when all committee members reach a unanimous consensus. The live chain differs from the archive chain in structure. Since the live chain only records the latest state of the models, smart contracts need to operate with models as transaction entries.
@@ V-A Transactional Models in the Dual-Chain Structure
  • $\bullet$
    $CH_{curr}\left(h\right)$: Considering model updating, it functions as the hash value for the current block. This role ensures that any changes to the model version resulting from updates, do not affect the validity of the current block.
  • $\bullet$
@@
  • $\bullet$
    versionversion: This field is updated with the corresponding version number for each update task, maintaining consistent versions in the global state.
There are also the common block body’s Merkle root $MR(h)$ and current block generation time timestamptimestamp in block header.
Block body. A block body contains multiple redactable transactions uniformly composed in a Merkle tree structure and stores model-related values. A redactable transaction contains the following fields, where we take transaction $Tx_{i,j}$ as an example, i.e., the $j$-th transaction in the body of block $i$.
  • $\bullet$
@@
The design of this block contributes to efficient and secure redactability in the blockchain. Due to the use of CH, transaction updates do not alter the block body’s Merkle root. However, the version field in the block header may cause collaborative updating issues with the block header. To ensure the consistency of transaction versions, we introduce the versionversion field in the block body. As a result, the validity of the link between blocks is maintained, and computational overhead is minimized.
Moreover, $CH\left(i,j\right)$ is a field in the block that needs verification for the block body. In this way, even if $URI_{w}\left(i,j\right)$ and $v(i,j)$ are updated, $CH(i,j)$ in the block body remains unchanged. Thus, $MR(h)$ in the block header also remains unchanged. These conditions satisfy the immutability of $CH_{curr}\left(h\right)$ and $CH(i,j)$. A valid block with an immutable block header can still be verified. On the other hand, this redactability does not compromise the tamper-resistance of the blockchain. The redactability of blocks is guaranteed by $CH_{sk}$ (trapdoor), an indispensable input for CH updates [10].
V-B Model Unlearning in Live Chain
@@
The parameter update equation related to $CH$, as specified in (6), ensures that the original $CH$ value remains valid after information updates. The committee inputs the original payload $\mathcal{T}(i,j)$, which includes the model $URI_{w}(i,j)$ before the update, along with the new payload $\mathcal{T}^{\prime}(i,j)$ containing the updated model $URI^{\prime}_{w}(i,j)$ and the unchanged reference list $R_{Tx}(i,j)$, the random number $v(i,j)$ before the update, and the corresponding private key $CH_{sk}$. It outputs the updated random number $v^{\prime}(i,j)$, as follows:
$\begin{split}CH\ update\left(\mathcal{T}(i,j),\mathcal{T}^{\prime}(i,j),v(i,j),CH_{sk}\right)\rightarrow v^{\prime}(i,j).\end{split}$ (6)
@@
After the transaction update is completed, the block header also needs to be updated. This is achieved by updating the version field, following a similar process to the transaction update. Once the committee completes the transaction and block updates via consensus, the information about which models have been updated and to what versions is broadcast throughout the live chain.
VI Blockchained Federated Unlearning Paradigms
@@
VI-A Paradigm 1 – Parallel Unlearning
Parallel unlearning refers to the committee members collectively reaching a single consensus for updating all to-be-unlearned models within an unlearning task. After updating all models and reaching a unified consensus, the results are recorded on the chain simultaneously. To illustrate the parallel unlearning paradigm, we employ the gradient ascent to demonstrate this process. We introduce a new gradient passing function that manipulates the ascending gradients for inherited models within the DAG using the ascending gradients from the starting models. This design ensures that updates to the starting model are efficiently propagated to all related inherited models, enabling parallel updates for single or multiple requests while reducing computational overhead.
@@ VI-A Paradigm 1 – Parallel Unlearning
$w^{\prime}_{y}=\sum_{p_{i}\in P}\frac{\alpha\nabla\theta_{s}}{\prod_{w_{j}\text{ on }p_{i}}N^{R_{j}}}+w_{y},$ (8)
where $P$ denotes the set of paths from $w_{s}$ to $w_{y}$, $P=\left\{p_{1},\ldots,p_{l}\right\}$, $w_{j}\text{ on }p_{i}$” denotes traversing each model node on path $p_{i}$ except $w_{s}$, $N^{R_{j}}$ denotes the number of models inherited by model node $w_{j}$ passing through path $p_{i}$, and $\alpha$ denotes discount factor. (8) handles model inheritance in FL, enabling the update of all relevant inherited models.
Update with multiple starting models. A user participating in a BlockFUL task contributes multiple models, and these models may have different gradients. During updates, these models serve as the starting models and have varying degrees of influence on the inherited models (i.e., the update of the inherited models may involve the computation of one or more starting model gradients). In other words, the inherited models may be affected by scenarios, e.g., the same gradient but different paths, or different gradients and paths.
@@
$w^{\prime}_{y}=\!\sum_{\nabla\theta^{k}_{s}\!\in\!\nabla\theta,P^{\nabla\theta^{k}_{s}}\!\in\!\mathbb{P}}\sum_{p^{\nabla\theta^{k}_{s}}_{i}\in P^{\nabla\theta^{k}_{s}}}\frac{\alpha\nabla\theta^{k}_{s}}{\prod_{w_{j}\text{ on }p_{i}}N^{R_{j}}}\!+w_{y},$ (9)
where $\nabla\theta=\left\{\nabla\theta^{1}_{s},\ldots,\nabla\theta^{h}_{s}\right\}$ is the set of multiple model gradients for a user, $p_{i}^{\nabla\theta^{k}_{s}}$ is the $i$-th path with gradient $\nabla\theta^{k}_{s}$, $P^{\nabla\theta^{k}_{s}}$ denotes the set of paths with gradient $\nabla\theta^{k}_{s}$, and $\mathbb{P}=\left\{P^{\nabla\theta^{1}_{s}},\ldots,P^{\nabla\theta^{h}_{s}}\right\}$ is the set of paths with all gradients.
@@
Next, we prove the boundedness of depths required for updating from the starting model to its inherited model. Let $\nabla\theta=\alpha\nabla\theta_{s}$, and consider $D=\{d_{1},d_{2},\ldots,d_{l}\}$ with $d_{i}$ being the depth of path $p_{i}$. We establish the following theorem.
@@
From (8), it readily follows that
@@
$\begin{split}&\sum_{p_{i}\in P}\frac{|\nabla\theta|}{\prod_{w_{j}\text{ on }p_{i}}N^{R_{j}}}\leqslant\sum_{d_{i}\in D}\frac{|\nabla\theta|}{2^{d_{i}}}\leqslant\varepsilon\Longrightarrow\sum_{d_{i}\in D}\frac{1}{2^{d_{i}}}\leqslant\frac{\varepsilon}{|\nabla\theta|}.\end{split}$ (16)
If $\frac{1}{\prod_{w_{j}\text{ on }p_{i}}N^{R_{j}}}=\frac{1}{2^{d_{i}}}$ at this point, $d_{i}$ can be obtained as
@@
According to (9), assuming model $w_{y}$ is influenced by multiple distinct gradient model updates, we can represent $B^{k}_{s}$ the gradient ascent required for $w_{y}$ under the influence of gradient $\nabla\theta^{k}_{s}$. Here, $B^{k}_{s}=\sum_{p^{\nabla\theta^{k}_{s}}_{i}\in P^{\nabla\theta^{k}_{s}}}\frac{1}{\prod\nolimits{{w_{j}\text{ on }p_{i}}}N^{R_{j}}}\times\alpha\nabla\theta^{k}_{s}$. Since multiple starting models are updated simultaneously, the gradient ascent process accumulates individual gradient ascent values. In other words, $B=B^{1}_{s}+B^{2}_{s}+\cdots+B^{h}_{s}$. By Theorem 1, the inherited models influenced by the starting models eventually stop updating when a certain depth is reached. As $\prod\nolimits_{w_{j}\text{ on }p_{i}}N^{R_{j}}$ increases, $B^{k}_{s}\leqslant\varepsilon$, resulting in $B\leqslant\varepsilon$. ∎
Parallel unlearning in BlockFUL. Users broadcast the updated model set $W^{\prime}$ and associated gradient messages to the network (Line 3, Algorithm 1), after which the algorithm iterates through $W^{\prime}$ using parallel execution loops (Line 4). The committee processes the updated models $W$ and their inherited models $N(W)$, applying gradient ascent updates (Lines 5-8), which are then executed on the blockchain. Next, the committee utilizes the reference list $R_{Tx}$ to establish reference relations for all transactions in the BlockFUL task and, using $CH_{sk}$, overwrites relevant transactions and versions on the blockchain (Lines 9-14) until all models in $W^{\prime}$ are processed. Finally, the committee verifies the validity of model updates (Lines 15-21), requiring only a single consensus operation for parallel unlearning to update both the FL model and blockchain. A unified hash value is computed from all prior updates, triggering an alarm and exiting the update block if verification fails; otherwise, the update is finalized, and the new models and blockchain data are officially recorded.
@@ 6  Send model_update message($w^{\prime}_{i}$);
8     $w^{\prime}_{j}$ $\leftarrow$ Committee.Update($w_{j}$, $\nabla\theta$, $\alpha$);
@@ 11    $\mathcal{T}^{\prime}(i,j)$, $v_{i,j},CH_{sk}$);
12     $R^{\prime}_{curr}$ $\leftarrow$ Committee.Overwrite(versionversion,
13    newnew versionversion, $R_{curr},CH_{sk}$);
14
@@
18
19while Consensus do
[…]
21  if $Hash(W^{\prime},N(W^{\prime}),v^{\prime}_{i,j},R^{\prime}_{curr})$ isis invalidinvalid then
@@
Algorithm 1 Parallel unlearning
VI-B Paradigm 2 – Sequential Unlearning
Sequential unlearning refers to a process where committee members execute a consensus process for each model update, i.e., a single DAG node, within an unlearning task. After updating one model and reaching a consensus, the results are recorded on the chain before moving on to the next model update. By updating models one by one and recording them on the chain, sequential unlearning ensures the compliance and procedural correctness of each operation. We employ the re-training method to demonstrate this process, with the same user model settings adopted as in gradient ascent. The re-training starts from the updated model $w^{\prime}_{s}$ and goes through all paths to the model $w_{y}$.
VI-B1 Re-aggregation
@@ VI-B2 Re-training
Sequential unlearning in BlockFUL. The client sequentially applies unlearning operations to each model in $W^{\prime}$ (Line 3, Algorithm 2). For the current model $w_{j}$, the client checks if it is the starting model. If $w_{j}\neq w_{i}^{\prime}$, the client locally performs re-aggregation (18) and re-training (19) (Lines 5–8), then broadcasts the updated model to the network (Line 9). The committee, using $CH_{sk}$ in the blockchain, executes the same update operations as in Algorithm 1 (Lines 10–12) and initiates the consensus verification process (Lines 13–20). Upon reaching consensus, the committee updates the blocks and parameters in IPFS and forwards model update messages to the next-hop clients to update the inherited models $N^{1}(W)$, $N^{2}(W)$, and so on, ensuring all inherited models in $N(W)$ are updated (Line 21).
@@
2Define All functions inherited from Algorithm 1;